# Covariant Quantum Error-Correcting Codes with Metrological Entanglement Advantage

### v1

@@ Acknowledgements.
$D_{P}(\Lambda,\Lambda+\mathcal{B})=O(\sqrt{d}n_{d}N^{(b-a)/2})~{}.$ (S21)
S5 Proof of Lemma 3
@@
Lemma 3: Consider the irrep of $SU(2)$ formed by $|J\!,\!M\rangle$, where $J=sN$, and the reduced density matrix $\rho_{M}:=\text{Tr}_{\bar{\alpha}}[|J\!,\!M\rangle\langle J\!,\!M|]$, where $\bar{\alpha}$ is the complement of the set of erased qudits $\alpha$. We then have the fidelity $f(\rho_{M},\rho_{M=0})=1-O(\frac{dM^{2}}{sN^{2}})=1-O(N^{2b+c-2})$ asymptotically for $M\sim N^{b}$, $d\sim N^{c}$, and $1>b\geq c$.
@@
×$\times\sum_{m_{1}=-j_{1}}^{j_{1}}\binom{2j_{1}}{j_{1}+m_{1}}\exp\left(\frac{Mm_{1}-m_{1}^{2}}{J}-\frac{Mm_{1}-m_{1}^{2}}{2J^{2}}(1-2j_{1})\right)$
@@
×$\times\sum_{m_{1}=-j_{1}}^{j_{1}}\binom{2j_{1}}{j_{1}+m_{1}}\left(1+\frac{Mm_{1}-m_{1}^{2}}{J}+\frac{(Mm_{1}-m_{1}^{2})^{2}}{2J^{2}}-\frac{(Mm_{1}-m_{1}^{2})}{2J^{2}}(1-2j_{1})+O(J^{-3})\right)$
@@
Theorem 2: Consider the code space $\mathfrak{C}=span\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J=sN$, $M_{\mathrm{max}}-M_{\mathrm{min}}\sim N^{b}$ and $b<1$. $\mathfrak{C}$ forms an $(\!(N,O((b-c)\log_{2}N))\!)$ AQECC against $d$-local erasures of size $d\sim N^{c}$, where $10\leq c\leq b<$ 1, with $\epsilon(d)=O(N^{b+c/2-1})$ if we take $\Delta\geq 2sd+1$.
@@
We consider the erasure errors at known locations expressed as $\mathcal{N}(\bullet)=\sum_{\alpha}p_{\alpha}|\alpha\rangle\langle\alpha|_{C}\otimes\mathcal{N}^{\alpha}(\bullet)$, where $\alpha$ denotes the set of qudits that are erased whose cardinal number is at most $d$, $C$ is the classical registry that records the set of lost qudits, and ${\cal N}^{\alpha}(\bullet):=\text{Tr}_{\alpha}(\bullet)$. To bound the code inaccuracy, following Bény and Oreshkov [18], we bound the entanglement fidelity between the complementary error channel $\widehat{N}(\rho)=(\Lambda+\mathcal{B})(\rho)$ and some channel $\Lambda(\rho)$, which needs to be independent of the input $\rho$. Note that the complementary channel of $\mathcal{N}^{\alpha}$ is $\widehat{\mathcal{N}}^{\alpha}(\bullet)=\text{Tr}_{\bar{\alpha}}(\bullet)$, where $\bar{\alpha}$ is the complement of $\alpha$. The complementary channel $\widehat{\mathcal{N}}$ to the full erasure channel is just the original channel, but with each $\mathcal{N}^{\alpha}\to\widehat{\mathcal{N}}^{\alpha}$ and with a different reference system $C^{\prime}$. We consider the channels $\Lambda(\rho):=\sum_{\alpha}p_{\alpha}|\alpha\rangle\langle\alpha|_{C^{\prime}}\otimes\mathcal{T}^{\alpha}(\rho)$ and $\mathcal{T}^{\alpha}(\rho):=\text{Tr}[\rho]\cdot\tau^{0}_{\alpha}$, both independent of the input $\rho$ and $\tau^{0}_{\alpha}:=\text{Tr}_{\bar{\alpha}}(|J,M\!=\!0\rangle\langle J,M\!=\!0|)$. The goal now is to bound the entanglement fidelity $F(\mathcal{\hat{N}}=\Lambda+\mathcal{B},\Lambda)$.
Recall that the code space is $\mathfrak{C}=span\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, and we use $|x\rangle_{L}=V^{\dagger}|J,M_{x}\rangle_{A}$, $x=1,\cdots,2^{k}$, to denote the basis in the logical space for $|J,M_{x}\rangle\in\mathfrak{C}$, where $V$ is the isometric encoding such that $VV^{\dagger}=\Pi$ is the projection onto the code space. Consider $\rho_{\alpha}^{x,x^{\prime}}=\text{Tr}_{\bar{\alpha}}[|J,M_{x}\rangle\langle J,M_{x^{\prime}}|]$, which is an at-most $d$-local operator. If we pick $\Delta\geq 2sd+1$, from Lemma 1, we should have $\rho_{\alpha}^{x,x^{\prime}}=0$ if $x\neq x^{\prime}$, assuming the size of $\alpha$ is at most $d$. We therefore will only consider the case $x=x^{\prime}$ and denote $\rho^{x}_{\alpha}:=\text{Tr}_{\bar{\alpha}}[|J,M_{x}\rangle\langle J,M_{x}|]$ for later convenience.
@@
To bound $f[(\widehat{\mathcal{N}}^{\alpha}\otimes\mathrm{id})(|\psi\rangle\langle\psi|),({\mathcal{T}}^{\alpha}\otimes\mathrm{id})(|\psi\rangle\langle\psi|)]$, recall that the auxiliary space of $[\widehat{\mathcal{N}}^{\alpha}\otimes\mathrm{id}]$ and $[{\mathcal{T}}^{\alpha}\otimes\mathrm{id}]$ is isomorphic to the input space, which is the logical space. For any state $|\psi\rangle=\sum_{x,y}A_{xy}|x\rangle\otimes|y\rangle$, consider the operators $A=\sum_{x,y}A_{xy}|x\rangle\langle y|$ and $B:=A^{T}=\sum_{x,y}A_{xy}|y\rangle\langle x|$ and the state $|\Phi\rangle=\sum_{x=1}^{2^{k}}|x\rangle\otimes|x\rangle$, we have $|\psi\rangle=A\otimes\mathrm{id}|\Phi\rangle=\mathrm{id}\otimes B|\Phi\rangle$. Therefore, we can write
@@
Since it holds for arbitrary $|\psi\rangle$, we have $F(\widehat{N},\Lambda)=1-O\left(\frac{dM_{\mathrm{max}}^{2}}{4sN^{2}}\right)$, or $\epsilon=D_{P}(\mathcal{\hat{N}},\Lambda)=O\left(\sqrt{\frac{d}{2s}}\frac{M_{\mathrm{max}}}{N}\right)$. In the scaling regime $M_{\mathrm{max}}-M_{\mathrm{min}}\sim N^{b}$ and $d\sim N^{c}$, we have $\epsilon=O(N^{b+c/2-1})$. The number of logical qubits is $k=\log_{2}\left(\frac{M_{\mathrm{max}}-M_{\mathrm{min}}}{\Delta}+1\right)$. Since we need to take $\Delta=\Omega(N^{c})$, we have $k=O((b-c)\log_{2}N)$ as claimed.
S7 Calculation of QFI loss and measurements

### v2

@@ Acknowledgements.
$D_{P}(\Lambda,\Lambda+\mathcal{B})=O(\sqrt{d}n_{d}N^{-(a-b)/2})~.$ (S22)
The above result illustrate the tradeoff between the parameters $d$, $n_{d}$, and the suppression factor $N^{-(a-b)/2}$
S5 Code Performance Against i.i.d. errors
While our Theorem 1 and Corollary 1 characterize the code’s performance against general errors, they assume knowledge of the error locations. This assumption effectively limits the number of Kraus operators. In particular, the code has a “room for error” of order $O(N^{-(a-b)/2})$. This means the code cannot effectively correct errors while achieving an asymptotically vanishing inaccuracy if the number of Kraus operators grows faster than this “room for error”, as indicated by Corollary 1. The assumption of knowing the locations of the error in Theorem 1 is therefore a way to restrict the number of Kraus operators of the error channel. However, we emphasize that, in the proofs of both Theorem 1 and Corollary 1, we assumed that each Kraus operator has an operator norm of order one, which is a worst-case assumption. On the other hand, one often expects that error processes involving more qudits occur with a suppressed probability, and thus the corresponding Kraus operators would have smaller operator norms. This consideration should allow for more Kraus operators, as long as the number of Kraus operators, weighted by their operator norms, grows more slowly than the “room for error.”
With this motivation, we analyze the code performance under the following i.i.d. error model. For simplicity, we consider the noise channel $\mathcal{N}=\bigotimes_{j=1}^{N}\mathcal{N}_{j}$, where each local noise channel acts as
$\mathcal{N}_{j}(\rho)=(1-p)\rho+pA_{j}\rho A_{j}^{\dagger}~,$ (S23)
with $A_{j}$ a unitary operator acting on site $j$, and $p$ the error probability.
This noise channel admits a Kraus representation
$\mathcal{N}(\rho)=\sum_{\vec{\ell}}E_{\vec{\ell}}\rho E_{\vec{\ell}}^{\dagger}~,$ (S24)
where $\vec{\ell}=(\ell_{1},\cdots,\ell_{N})\in\{0,1\}^{N}$ is a binary string, and we use $\ell$ to denote its Hamming weight, defined as the number of entries equal to $1$ in $\vec{\ell}$.
The corresponding Kraus operator takes the form
$E_{\vec{\ell}}:=\sqrt{p_{\ell}}E_{\ell_{1}}E_{\ell_{2}}\cdots E_{\ell_{N}}~,$ (S25)
with $p_{\ell}=p^{\ell}(1-p)^{N-\ell}$, where $E_{\ell_{j}=0}=I_{j}$ is the identity operator on site $j$, and $E_{\ell_{j}=1}=A_{j}$ is the error operator on site $j$.
It is useful to decompose the channel into contributions from errors acting on at most $d$ sites, and those acting on more than $d$ sites:
$\mathcal{N}=P_{d}\mathcal{N}_{d}+P_{>d}\mathcal{N}_{>d}~,$ (S26)
where $P_{d}=\sum_{\ell=0}^{d}\binom{N}{\ell}p^{\ell}(1-p)^{N-\ell}$ and $P_{>d}=1-P_{d}=\sum_{\ell=d+1}^{N}\binom{N}{\ell}p^{\ell}(1-p)^{N-\ell}$ are the respective probabilities. The maps $\mathcal{N}_{d}$ and $\mathcal{N}_{>d}$ are completely positive and trace-preserving (CPTP), describing the $d$-qudit errors and the remaining higher-weight errors, respectively. More specifically, $\mathcal{N}_{d}(\rho)=\sum_{\vec{\ell}:\ell\leq d}K_{\vec{\ell}}\rho K_{\vec{\ell}}^{\dagger}$, where $K_{\vec{\ell}}=E_{\vec{\ell}}/\sqrt{P_{d}}$. For later convenience, we set $K_{\vec{\ell}}=0$ if $\ell>d$. As we will show, our code can protect against this noise model by correcting $d$-qudit errors $\mathcal{N}_{d}$, with $d\leq(c+o(1))\log N$, provided the error probability $p$ satisfies $p\leq(c_{0}^{2}+o(1))N^{-2}\log^{2}N$. We now state this result.
Corollary 2.
Consider the code space $\mathfrak{C}=span\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J\sim N^{a}$, $M_{\mathrm{max}}\!-\!M_{\mathrm{min}}\sim N^{b}$, $b<a\leq 1$. Under the i.i.d. noise channel $\mathcal{N}=\bigotimes_{j=1}^{N}\mathcal{N}_{j}$, where $\mathcal{N}_{j}(\rho)=(1-p)\rho+pA_{j}\rho A_{j}^{\dagger}$ for a unitary matrix $A_{j}$ on site $j$, $\mathfrak{C}$ forms an $(\!(N,O(b\log_{2}N))\!)$ AQECC by correcting $d$-qudit errors with $d=(c+o(1))\log N$, achieving a code inaccuracy $\epsilon=O(\sqrt{\log N}N^{-\beta/2})$, if $p=(c_{0}^{2}+o(1))(N^{-2}\log^{2}N)$, $\Delta\geq 4sd+1$, and $\beta=a-b-c_{0}-2c>0$.
Proof.
We aim to lower bound the entanglement fidelity $\max_{\mathcal{R}}F(\mathcal{R}\mathcal{N},\mathrm{id})$, in order to demonstrate that the purified distance is small. Note that, using the concavity of the fidelity, we have
$f(\mathcal{R}\mathcal{N}\otimes\mathrm{id}(|\psi\rangle\langle\psi))\geq P_{d}f(\mathcal{R}\mathcal{N}_{d}\otimes\mathrm{id}(|\psi\rangle\langle\psi))+P_{>d}f(\mathcal{R}\mathcal{N}_{>d}\otimes\mathrm{id}(|\psi\rangle\langle\psi))$ (S27)
for any state $|\psi\rangle$ and channel $\mathcal{R}$. Suppose $|\psi^{*}\rangle$ minimizes $f(\mathcal{R}\mathcal{N}\otimes\mathrm{id}(|\psi\rangle\langle\psi|))$. We then have $f(\mathcal{R}\mathcal{N}_{d}\otimes\mathrm{id}(|\psi^{*}\rangle\langle\psi^{*}|))\geq\min_{\psi}f(\mathcal{R}\mathcal{N}_{d}\otimes\mathrm{id}(|\psi\rangle\langle\psi|))=F(\mathcal{R}\mathcal{N}_{d},\mathrm{id})$, and similarly $f(\mathcal{R}\mathcal{N}_{>d}\otimes\mathrm{id}(|\psi^{*}\rangle\langle\psi^{*}|))\geq F(\mathcal{R}\mathcal{N}_{>d},\mathrm{id})$. We therefore have
$F(\mathcal{R}\mathcal{N},\mathrm{id})\geq P_{d}F(\mathcal{R}\mathcal{N}_{d},\mathrm{id})+P_{>d}F(\mathcal{R}\mathcal{N}_{>d},\mathrm{id})$ (S28)
for any quantum channel $\mathcal{R}$.
Suppose we can lower bound $\max_{\mathcal{S}}F(\mathcal{S}\mathcal{N}_{d},\mathrm{id})\geq 1-\epsilon_{d}$, and assume $\mathcal{S}^{*}$ is the channel that minimizes the left-hand side. Then we have
$\max_{\mathcal{R}}F(\mathcal{R}\mathcal{N},\mathrm{id})$ $\geq F(\mathcal{S}^{*}\mathcal{N},\mathrm{id})$
$\geq P_{d}F(\mathcal{S}^{*}\mathcal{N}_{d},\mathrm{id})+P_{>d}F(\mathcal{S}^{*}\mathcal{N}_{>d},\mathrm{id})$
$\geq P_{d}(1-\epsilon_{d})=1-(P_{>d}+P_{d}\epsilon_{d})~,$ (S29)
where we have used the concavity of entanglement fidelity Eq. (S28), $F(\mathcal{S}^{*}\mathcal{N}_{d},\mathrm{id})\geq(1-\epsilon_{d})$, and $F(\mathcal{S}^{*}\mathcal{N}_{>d},\mathrm{id})\geq 0$. The purified distance is therefore
$\min_{\mathcal{R}}D_{P}(\mathcal{R}\mathcal{N},\mathrm{id})\leq\sqrt{1-(1-P_{>d}-P_{d}\epsilon_{d})^{2}}\leq\sqrt{2(P_{>d}+P_{d}\epsilon_{d})}~.$ (S30)
Physically, this inequality indicates that the code inaccuracy can be made small as long as the probability $P_{>d}$ is small, and the code achieves small inaccuracy $\epsilon_{d}$ when correcting $d$-local errors $\mathcal{N}_{d}$. As we will show below, we can obtain an upper bound for $\min_{\mathcal{R}}D_{P}(\mathcal{R}\mathcal{N},\mathrm{id})$ by bounding $P_{>d}$ and $P_{d}\epsilon_{d}$ from above.
First, we lower bound $\max_{\mathcal{S}}F(\mathcal{S}\mathcal{N}_{d},\mathrm{id})$. We follow the same strategy as the one used in the proofs of Theorem 1 and Corollary 1 by examining Knill-Laflamme conditions for the noise channel $\mathcal{N}_{d}(\rho)=\sum_{\vec{\ell}}K_{\vec{\ell}}\rho K_{\vec{\ell}}^{\dagger}$, where, for convenience, we set $K_{\vec{\ell}}=0$ if $\ell>d$. Considering the code space $\mathfrak{C}=\text{span}\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, from Lemma 1, we have $\langle J\!,\!m|K_{\vec{r}}^{\dagger}K_{\vec{t}}|J\!,\!n\rangle=0$ if $|m-n|\geq 4sd+1$, which is satisfied if $\Delta\geq 4sd+1$. For the diagonal Knill-Laflamme conditions, using Lemma 2, we have
$|\langle J\!,\!m|K^{\dagger}_{\vec{r}}K_{\vec{t}}|J\!,\!m\rangle-\langle J\!,\!M_{\min}|K^{\dagger}_{\vec{r}}K_{\vec{t}}|J\!,\!M_{\min}\rangle|=\frac{\sqrt{p_{r}p_{t}}}{P_{d}}\cdot O(dN^{-(a-b)})~,$ (S31)
where we have used $\|K_{\vec{\ell}}\|_{\text{op}}=\sqrt{p_{\ell}}$, where $p_{\ell}=(1-p)^{N-\ell}p^{\ell}$ and $\ell$ is the Hamming weight of the binary vector $\vec{\ell}$.
Similar to the proof of Theorem 1, we take $\lambda_{\vec{r}\vec{t}}=\langle J\!,\!M_{\min}|K^{\dagger}_{\vec{r}}K_{\vec{t}}|J\!,\!M_{\min}\rangle$, and the channels $\Lambda$ and $\hat{\mathcal{N}}_{d}=\Lambda+\mathcal{B}$ to be $\Lambda(\rho)=\sum_{\vec{r},\vec{t}}\lambda_{\vec{r}\vec{t}}\mathrm{Tr}[\rho]|\vec{r}\rangle\langle\vec{t}|$ and $(\Lambda+\mathcal{B})(\rho)=\sum_{\vec{r},\vec{t}}\mathrm{Tr}[K_{\vec{t}}\rho K^{\dagger}_{\vec{r}}]|\vec{r}\rangle\langle\vec{t}|$, respectively. To bound the entanglement fidelity $F(\hat{\mathcal{N}}_{d},\Lambda)$, consider the state $|\Phi\rangle=\sum_{x=1}^{2^{k}}|x\rangle\otimes|x\rangle$. We then have
[…]
where $\Gamma^{x}_{\vec{r}\vec{t}}=\langle J\!,\!M_{x}|K^{\dagger}_{\vec{r}}K_{\vec{t}}|J\!,\!M_{x}\rangle$. Analogous to Eq. (S4), we have
$f((\Lambda+\mathcal{B})\otimes\mathrm{id}(|\psi\rangle\langle\psi|),\Lambda\otimes\mathrm{id}(|\psi\rangle\langle\psi|))\geq\sum_{x=1}^{2^{k}}\langle x|B^{\dagger}B|x\rangle\cdot f\left(\sum_{\vec{r},\vec{t}}\Gamma_{\vec{r}\vec{t}}^{x}|\vec{r}\rangle\langle\vec{t}|,\sum_{\vec{r},\vec{t}}\lambda_{\vec{r}\vec{t}}|\vec{r}\rangle\langle\vec{t}|\right)$ (S34)
due to the joint concavity of fidelity. We then proceed to bound the fidelity between the density matrices $\Gamma_{x}:=\sum_{\vec{r},\vec{t}}\Gamma_{\vec{r}\vec{t}}^{x}|\vec{r}\rangle\langle\vec{t}|$ and $\lambda_{0}:=\sum_{\vec{r},\vec{t}}\lambda_{\vec{r}\vec{t}}|\vec{r}\rangle\langle\vec{t}|$ using the trace distance, obtaining
$\|\Gamma_{x}-\lambda_{0}\|_{1}$ $\leq\sum_{\vec{r},\vec{t}}|\Gamma_{\vec{r}\vec{t}}^{x}-\lambda_{\vec{r}\vec{t}}|=\sum_{r=0}^{d}\sum_{t=0}^{d}\binom{N}{r}\binom{N}{t}\frac{1}{P_{d}}(1-p)^{N-\frac{1}{2}(r+t)}p^{\frac{1}{2}(r+t)}\cdot O(dN^{-(a-b)})$
$=P_{d}^{-1}(1-p)^{N}\mathbf{F}_{d}^{2}\cdot O(dN^{-(a-b)})~,$ (S35)
where
$\mathbf{F}_{d}:=\sum_{\ell=0}^{d}\binom{N}{\ell}\left(\frac{p}{1-p}\right)^{\frac{\ell}{2}}=\sum_{\ell=0}^{d}\binom{N}{\ell}p_{1}^{\ell}=\sum_{\ell=0}^{d}\binom{N}{\ell}p_{2}^{\ell}(1-p_{2})^{N-\ell}\cdot(1+p_{1})^{N}~,$ (S36)
where $p_{1}:=\sqrt{p/(1-p)}$ and $p_{2}:=p_{1}/(1+p_{1})$. We are interested in the regime $d=(c+o(1))\log N$. Since we assume $p=(c_{0}^{2}+o(1))N^{-2}\log^{2}N$, we have $p_{2}=\sqrt{p}/(\sqrt{1-p}+\sqrt{p})=(c_{0}+o(1))N^{-1}\log N$, or $Np_{2}=(c_{0}+o(1))\log N$. Assuming $d=(1-\delta)Np_{2}$ for some $\delta\in(0,1)$, the Chernoff bound for the binomial distribution gives
$\sum_{\ell=0}^{d}\binom{N}{\ell}p_{2}^{\ell}(1-p_{2})^{N-\ell}\leq\exp(-\frac{1}{2}\delta^{2}Np_{2})=O(N^{-\frac{1}{2}c_{0}\delta^{2}})~.$ (S37)
Since $p_{1}=p_{2}/(1-p_{2})$, we have
$(1+p_{1})^{N}=\exp(N\log(1+p_{1}))=\exp(-N\log(1-p_{2}))=O(\exp(Np_{2}))=O(N^{c_{0}})~.$ (S38)
Similarly, we have
$(1-p)^{N}=\exp(N\log(1-p))=O(\exp(-Np))=O(\exp(-c_{0}^{2}N^{-1}\log^{2}N))~,$ (S39)
which is sub-leading compared to $N^{\alpha}$ for any exponent $\alpha$. We therefore have
$\|\Gamma_{x}-\lambda_{0}\|_{1}=P_{d}^{-1}\cdot O(dN^{-\alpha})~,$ (S40)
where $\alpha=a-b+c_{0}(\delta^{2}-2)=a-b-c_{0}-2c+c^{2}/c_{0}+o(1)$, since $\delta=1-d/(Np_{2})=1-c/c_{0}+o(1)$.
Using Fuchs-van de Graaf inequality Eq. (S1), we have $f(\Gamma_{x},\lambda_{0})\geq 1-\frac{1}{2}\|\Gamma_{x}-\lambda_{0}\|$, and therefore
$\max_{\mathcal{S}}F(\mathcal{S}\mathcal{N}_{d},\mathrm{id})=\max_{\mathcal{S}^{\prime}}F(\hat{\mathcal{N}}_{d},\mathcal{S}^{\prime}\circ\text{Tr})\geq F(\hat{\mathcal{N}}_{d},\Lambda)\geq 1-P_{d}^{-1}\cdot O(dN^{-\alpha})~,$ (S41)
where we used the result in Ref. [18] for the first equality and the fact that $\text{Tr}[\bullet]$ is the complimentary channel of the identity channel $\mathrm{id}$.
Finally, to bound $P_{>d}=\sum_{\ell=d+1}^{N}\binom{N}{\ell}p^{\ell}(1-p)^{N-\ell}$, we have
$P_{>d}\leq\binom{N}{d+1}p^{d+1}<\left(\frac{Nep}{(d+1)}\right)^{d+1}<\left(\frac{Nep}{d}\right)^{d+1}=O(\exp(d\log\frac{Nep}{d}))=O(N^{-c\log N})~,$ (S42)
which approaches zero faster than $N^{-\alpha}$ for any $\alpha>0$, and we have used $\binom{N}{m}<(\frac{Ne}{m})^{m}$.
Substituting the results from Eq. (S41) and Eq. (S42) into Eq. (S30), we have
$\epsilon=\min_{\mathcal{R}}D_{P}(\mathcal{R}\mathcal{N},\mathrm{id})=O(\sqrt{\log N}N^{-\beta/2})~,$ (S43)
where $\beta:=a-b-c_{0}-2c<\alpha$. We therefore obtain the result of Corollary 2 as long as $\beta=a-b-c_{0}-2c>0$.
S6 Proof of Lemma 3
@@
Lemma 3: Consider the irrep of SU(2) formed by $|J\!,\!M\rangle$, where $J=sN$, and the reduced density matrix $\rho_{M}:=\mathrm{Tr}_{\bar{\alpha}}[|J\!,\!M\rangle\langle J\!,\!M|]$, where $\bar{\alpha}$ is the complement of the set of erased qudits $\alpha$. We have the fidelity $f(\rho_{M},\rho_{M=0})=1-O(\frac{dM^{2}}{sN^{2}})=1-O(N^{2b+c-2})$ asymptotically for $M\sim N^{b}$, $d\sim N^{c}$, and $1>b\geq c$.
@@
$\times\sum_{m_{1}=-j_{1}}^{j_{1}}\binom{2j_{1}}{j_{1}+m_{1}}\exp\left(\frac{Mm_{1}-m_{1}^{2}}{J}-\frac{Mm_{1}-m_{1}^{2}}{2J^{2}}(1-2j_{1})\right)$
@@
$\times\sum_{m_{1}=-j_{1}}^{j_{1}}\binom{2j_{1}}{j_{1}+m_{1}}\left(1+\frac{Mm_{1}-m_{1}^{2}}{J}+\frac{(Mm_{1}-m_{1}^{2})^{2}}{2J^{2}}-\frac{(Mm_{1}-m_{1}^{2})}{2J^{2}}(1-2j_{1})+O(J^{-3})\right)$
@@
Theorem 2: Consider the code space $\mathfrak{C}=\text{span}\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J=sN$, $M_{\mathrm{max}}-M_{\mathrm{min}}\sim N^{b}$ and $b<1$. $\mathfrak{C}$ forms an $(\!(N,O[(b\!-\!c)\log_{2}N])\!)$ AQECC against heralded $d$-local erasures of size $d\sim N^{c}$, where $0\leq c\leq b<1$, with $\epsilon(d)=O(N^{-(1-b-c/2)})$ if we take $\Delta\geq 2sd+1$.
@@
We consider the erasure errors at known locations expressed as $\mathcal{N}(\bullet)=\sum_{\alpha}p_{\alpha}|\alpha\rangle\langle\alpha|_{C}\otimes\mathcal{N}^{\alpha}(\bullet)$, where $\alpha$ denotes the set of qudits that are erased whose cardinal number is at most $d$, $C$ is the classical registry that records the set of lost qudits, and ${\cal N}^{\alpha}(\bullet):=\mathrm{Tr}_{\alpha}(\bullet)$. To bound the code inaccuracy, following Bény and Oreshkov [18], we bound the entanglement fidelity between the complementary error channel $\widehat{N}(\rho)=(\Lambda+\mathcal{B})(\rho)$ and some channel $\Lambda(\rho)$, which needs to be independent of the input $\rho$. Note that the complementary channel of $\mathcal{N}^{\alpha}$ is $\widehat{\mathcal{N}}^{\alpha}(\bullet)=\mathrm{Tr}_{\bar{\alpha}}(\bullet)$, where $\bar{\alpha}$ is the complement of $\alpha$. The complementary channel $\widehat{\mathcal{N}}$ to the full erasure channel is just the original channel, but with each $\mathcal{N}^{\alpha}\to\widehat{\mathcal{N}}^{\alpha}$ and with a different reference system $C^{\prime}$. We consider the channels $\Lambda(\rho):=\sum_{\alpha}p_{\alpha}|\alpha\rangle\langle\alpha|_{C^{\prime}}\otimes\mathcal{T}^{\alpha}(\rho)$ and $\mathcal{T}^{\alpha}(\rho):=\mathrm{Tr}[\rho]\cdot\tau^{0}_{\alpha}$, both independent of the input $\rho$ and $\tau^{0}_{\alpha}:=\mathrm{Tr}_{\bar{\alpha}}(|J,M\!=\!0\rangle\langle J,M\!=\!0|)$. The goal now is to bound the entanglement fidelity $F(\mathcal{\hat{N}}=\Lambda+\mathcal{B},\Lambda)$.
Recall that the code space is $\mathfrak{C}=span\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, and we use $|x\rangle_{L}=V^{\dagger}|J,M_{x}\rangle_{A}$, $x=1,\cdots,2^{k}$, to denote the basis in the logical space for $|J,M_{x}\rangle\in\mathfrak{C}$, where $V$ is the isometric encoding such that $VV^{\dagger}=\Pi$ is the projection onto the code space. Consider $\rho_{\alpha}^{x,x^{\prime}}=\mathrm{Tr}_{\bar{\alpha}}[|J,M_{x}\rangle\langle J,M_{x^{\prime}}|]$, which is an at-most $d$-local operator. If we pick $\Delta\geq 2sd+1$, from Lemma 1, we should have $\rho_{\alpha}^{x,x^{\prime}}=0$ if $x\neq x^{\prime}$, assuming the size of $\alpha$ is at most $d$. We therefore will only consider the case $x=x^{\prime}$ and denote $\rho^{x}_{\alpha}:=\mathrm{Tr}_{\bar{\alpha}}[|J,M_{x}\rangle\langle J,M_{x}|]$ for later convenience.
@@
To bound $f[(\widehat{\mathcal{N}}^{\alpha}\otimes\mathrm{id})(|\psi\rangle\langle\psi|),({\mathcal{T}}^{\alpha}\otimes\mathrm{id})(|\psi\rangle\langle\psi|)]$, recall that the auxiliary space of $\widehat{\mathcal{N}}^{\alpha}\otimes\mathrm{id}]$ and ${\mathcal{T}}^{\alpha}\otimes\mathrm{id}]$ is isomorphic to the input space, which is the logical space. For any state $|\psi\rangle=\sum_{x,y}A_{xy}|x\rangle\otimes|y\rangle$, consider the operators $A=\sum_{x,y}A_{xy}|x\rangle\langle y|$ and $B:=A^{T}=\sum_{x,y}A_{xy}|y\rangle\langle x|$ and the state $|\Phi\rangle=\sum_{x=1}^{2^{k}}|x\rangle\otimes|x\rangle$, we have $|\psi\rangle=A\otimes\mathrm{id}|\Phi\rangle=\mathrm{id}\otimes B|\Phi\rangle$. Therefore, we can write
@@
Since it holds for arbitrary $|\psi\rangle$, we have $F(\widehat{N},\Lambda)=1-O\left(\frac{dM_{\mathrm{max}}^{2}}{4sN^{2}}\right)$, or $\epsilon=D_{P}(\mathcal{\hat{N}},\Lambda)=O\left(\sqrt{\frac{d}{2s}}\frac{M_{\mathrm{max}}}{N}\right)$. In the scaling regime $M_{\mathrm{max}}-M_{\mathrm{min}}\sim N^{b}$ and $d\sim N^{c}$, we have $\epsilon=O(N^{b+c/2-1})$. The number of logical qubits is $k=\log_{2}\left(\frac{M_{\mathrm{max}}-M_{\mathrm{min}}}{\Delta}+1\right)$. Since $d\sim N^{c}$, we need to take the spacing of the magnetic quantum number $\Delta=\Omega(N^{c})$. We therefore have $k=O((b-c)\log_{2}N)$ as claimed.
S8 Calculation of QFI loss and measurements